10. (10 points) Sodium azide, NaNs, is used to provide gas to inflate automobile air bags. What mass of sodium azide is needed to provide the nitrogen needed to inflate a 70.0 L bag to a pressure of 1.3 atm at 25°C? 2 NaN3 (s) → 2 Na (s) + 3 N2(g)
Answer #1

we have:

P = 1.3 atm

V = 70.0 L

T = 25.0 oC

= (25.0+273) K

= 298 K

find number of moles using:

P * V = n*R*T

1.3 atm * 70 L = n * 0.08206 atm.L/mol.K * 298 K

n = 3.719 mol

This is number of moles of N2 required

Now from reaction,

mol of NaN3 required = (2/3)*moles of N2

=(2/3)*3.719 mol

= 2.479 mol

Molar mass of NaN3 = 1*MM(Na) + 3*MM(N)

= 1*22.99 + 3*14.01

= 65.02 g/mol

we have below equation to be used:

mass of NaN3,

m = number of mol * molar mass

= 2.479 mol * 65.02 g/mol

= 161 g

Answer: 161 g

Feel free to comment below if you have any doubts or if this answer do not work